Problem 3: A study examines chocolates effects on blood vessel function in healthy people. In the randomized, double-blind, show work please
Answer #1

a)

Sample #1   ---->   1              
mean of sample 1,    x̅1=   1.300              
standard deviation of sample 1,   s1 =    2.320              
size of sample 1,    n1=   11              
                      
Sample #2   ---->   2              
mean of sample 2,    x̅2=   -0.960              
standard deviation of sample 2,   s2 =    1.580              
size of sample 2,    n2=   10              
                      
difference in sample means =    x̅1-x̅2 =    1.3000   -   -1.0   =   2.26

DF = min(n1-1 , n2-1 )=   9              
t-critical value , t* =    2.2622   (excel formula =t.inv(α/2,df)          
                  

                  
                  
std error , SE =    √(s1²/n1+s2²/n2) =    0.860          
margin of error, E = t*SE =    2.2622   *   0.860   =   1.94
                  
difference of means = x̅1-x̅2 =    1.3000   -   -0.960   =   2.260


confidence interval is                   
Interval Lower Limit = (x̅1-x̅2) - E =    2.2600   -   1.9446   =   0.3154
Interval Upper Limit = (x̅1-x̅2) + E =    2.2600   -   1.9446   =   4.2046

CI ( 0.32 , 4.20)

...............

b)

one can be 95% confident that true mean difference between groups will lie between these limits

.................

Please let me know in case of any doubt.

Thanks in advance!


Please upvote!